# The Riemann Hypothesis is false

### v46

@@
ABSTRACT. Let $\Theta$ denote the supremum of the real parts of the zeros of the Riemann zeta function. We demonstrate that $\Theta=1$, which entails the existence of infinitely many Riemann zeros off the critical line (thus disproving the Riemann Hypothesis (RH)). The paper is concluded by a brief discussion of why our argument doesn’t work for both Weil and Beurling zeta functions whose analogues of the RH are known to be true.
@@
 [^1] Data availability statement: This manuscript has no associated data. Introduction. The Riemann zeta function is a function of the complex variable $s$, defined in the half-plane $\Re(s)>1$ by $\zeta(s):=\sum_{n=1}^{\infty}n^{-s}$ and in the whole complex plane by holomorphic continuation. Euler noticed that for $\Re(s)>1$, $\zeta(s)$ can be expressed as a product $\prod_{p}(1-p^{-s})^{-1}$ over the entire set of primes, which entails that $\zeta(s)\neq 0$ for $\Re(s)>1$. It can be shown that $\zeta(s)$ extends to $\mathbb{C}$ as a meromorphic function with only a simple pole at $s=1$, with residue $1$, and satisfies the functional equation $\xi(s)=\xi(1-s)$, where $\xi(z)=\frac{1}{2}z(z-1)\pi^{-z/2}\Gamma(\frac{1}{2}z)\zeta(z)$ and $\Gamma(w)=\int_{0}^{\infty}e^{-x}x^{w-1}\mathrm{d}x$. Let $\chi$ be a primitive Dirichlet character and $L(s,\chi)$ be the corresponding L-function. Most of the analytic properties for $\zeta$ are also known to hold for $L(s,\chi)$ [5, Chapter 10]. Let $\Theta_{\chi}$ be the supremum of the real parts of the zeros of $L(s,\chi)$. The Generalised Riemann Hypothesis (GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$. For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta,\sigma=\Re(s)$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
Lemma 1. (Theorem 4.11 of [7]). We have $\zeta(s)=\sum_{n\leq x}n^{-s}-\frac{x^{1-s}}{1-s}+O(x^{-\sigma})$ uniformly for $\sigma\geq\sigma_{0}>0$ and $|\Im(s)|<2\pi x/C$, where $C$ is a given constant $>1$.
@@
Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
@@
uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
@@
where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant, $\chi(s)=\frac{\zeta(s)}{\zeta(1-s)}$ and $|\chi(\sigma-it)|=|\chi(\sigma+it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78]. Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1+it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
@@
for $-1\leq\sigma\leq 2.$ Combining (1) and (6) gives $\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{8}T.$ By the log-derivative of the functional equation for $\zeta$, we have
@@
We now treat the ranges $m+1\leq 2n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq 2n\leq m-1$, where $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. For $m+1\leq 2n\leq m(1+\varepsilon)$, put $2n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(2n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
@@
$=\frac{\Lambda(m)}{em^{iT}}\sum_{0<r\leq\varepsilon m-1}\Big{(}\frac{(m+r)^{iT}}{r}+O(1/m)\Big{)}$ (16)
@@
Let $J(u)=\sum_{0<r\leq u}e^{2\pi if(r)}=\sum_{0<r\leq M}e^{2\pi if(r)}$ where $u\leq m\varepsilon,M=M(u)$ is the largest integer $\leq u$, $f(r)=\frac{T\log(1+\frac{r}{m})}{2\pi}$ and $m\in\mathbb{N}$. Then $f^{\prime}(r)=\frac{T}{2\pi(r+m)}$. Put $\alpha=f^{\prime}(M),\beta=f^{\prime}(0)$ and note that $\beta-\alpha\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$ uniformly for $0\leq u\leq m\varepsilon$. Thus if $\eta$ is any constant $\in(0,1)$, then the range $[\alpha-\eta,\beta+\eta]$ contains at most two integers. By Lemma 4.2 of Titchmarsh [7], we have $\Big{|}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x\Big{|}\leq\frac{4}{f^{\prime}(M)}\ll 1.$ Thus $\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x\ll 1$ and by Lemma 4.7 of the same book we have
$J(u)=\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x+O(1)\ll 1$ (18)
uniformly for $1\leq u\leq m\varepsilon.$ Then
$\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}$ $=\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{2\pi if(r)}$ (19)
[…]
Since $J(u)\ll 1$ uniformly for $1\leq u\leq m\varepsilon,$ it follows from (21) that
$\sum_{T^{1/4}<r\leq m\varepsilon-1}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}\ll T^{-1/4}=\varepsilon$. Applying this into (17) yields
$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log m).$ (22)
Putting $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O(\frac{r^{2}\log^{4}T}{m})}=1+O(\frac{r^{2}\log^{4}T}{m})$ into the right-hand side of (22) gives
$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{6}T).$ (23)
Inserting (23) into (17) and summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$, yields
@@
$+O(T^{3/4}\log^{6}T).$ (25)
For $\frac{m}{1+\varepsilon}\leq 2n\leq m-1,$ put $2n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ so that $\log(2n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$. Thus by arguing as above, one arrives at
@@
$=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log^{6}T).$ (29)
Let $\{{y\}}$ denote the fractional part of $y$ and $\mid\mid y\mid\mid$ denote the distance to the nearest integer to $y$. Note that if $u>0$ is not an integer, we have
$\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi\Big{(}\frac{1}{2}-\{{u\}}+O\Big{(}\frac{1}{1+X||u||}\Big{)}\Big{)}$ (30)
as $X\rightarrow\infty$. 222see e.g. https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O\Big{(}\frac{1}{1+T^{1/4}||T/2\pi m||}\Big{)}\Big{)}.$ (31)
Let $U_{m},V_{m}\in\mathbb{N},T=\frac{U_{m}}{V_{m}}$ and $m\mid U_{m}$. By Dirichlet’s theorem on rational approximations to irrationals, note that $\frac{1}{2\pi}$ can be approximated by infinitely many irreducible rational numbers $\frac{p_{m}}{q_{m}}$ with error $O({q_{m}}^{-2})$. Let: $V_{m}\mid p_{m}$, $\gcd(q_{m},U_{m})=1,\frac{U_{m}}{mV_{m}{q_{m}}^{2}}=o(1/q_{m})$ and $q_{m}\ll(\log T)^{6}$ uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. Note that: $\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}$ is irreducible, $\Big{|}\Big{|}\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}\Big{|}\Big{|}\gg\frac{1}{q_{m}}$ and $\frac{T}{2\pi m}=\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}+O\Big{(}\frac{U_{m}}{mV_{m}{q_{m}}^{2}}\Big{)}=\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}+o(1/q_{m})$. Hence $||T/2\pi m||\gg\frac{1}{q_{m}}\gg(\log T)^{-6}$ uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. Inserting (31) into (29) and applying the fact that $||T/2\pi m||\gg(\log T)^{-6}$, gives
@@
$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}\Big{)}+O(T^{3/4}\log^{6}T).$ (33)
@@
since $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.187]. Plugging (38) into (33) gives
@@
Remark 2. Our argument crucially relies on:
(i) The functional equation for $\zeta$, and
(ii) Theorem 4.11 of [7].
Both (i) and (ii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers. Hence our approach wouldn’t work for such, whose analogues of the RH are known to be true. Also, (ii) does not hold for the Weil zeta functions, which are known to satisfy the RH.
@@
Remark 4. The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need our Lemma 1 to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log t)^{-1}$. However, this is not possible (as can be seen e.g by estimating $\zeta(\sigma+it)=\sum_{n\leq\frac{T}{2}}n^{-\sigma-it}=\sum_{n\leq T/2}n^{(-\sigma+\sigma_{t})-\sigma_{t}-it}$ for any fixed $\sigma>0$ via Abel summation, under the assumption that $\sum_{n\leq T/2}n^{-\sigma_{t}-it}=\zeta(\sigma_{t}+it)+O(1/\log t)$). It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}=o(1)$. Indeed, he begins his argument by quoting the Euler-McLaurin summation formula for $\zeta(\sigma+it)$ (equation 3.5.3). This is initially defined for $\sigma>1$, and then is extended by holomorphic continuation to $\sigma>0$. Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh assumes that $\sigma$ is bounded below by some constant $\sigma_{0}>0$. For the same reason, one cannot apply any of the best known zero-free regions of $\zeta$ (e.g. Vinogradov-Koborov) to bound $\Re(\rho_{T})$ in (54).

### v47

@@
ABSTRACT. Let $\Theta$ denote the supremum of the real parts of the zeros of the Riemann zeta function $\zeta(s)$. By carefully exploiting the Landau-Gonek explicit formula, the Dirichlet polynomial approximation of $\zeta(s)$ and other classical results in analytic number theory, we rigorously demonstrate that $\Theta=1$. This entails the existence of infinitely many Riemann zeros off the critical line (thus disproving the Riemann Hypothesis (RH)). The paper is concluded by a brief discussion of why our argument does not work for both Weil and Beurling zeta functions, whose analogues of the RH are known to be true.
@@
 [^1] Data availability statement: This manuscript has no associated data. Introduction. The Riemann zeta function is a function of the complex variable $s$, defined in the half-plane $\Re(s)>1$ by $\zeta(s):=\sum_{n=1}^{\infty}n^{-s}$ and in the whole complex plane by holomorphic continuation. Euler noticed that for $\Re(s)>1$, $\zeta(s)$ can be expressed as a product $\prod_{p}(1-p^{-s})^{-1}$ over the entire set of primes, which entails that $\zeta(s)\neq 0$ for $\Re(s)>1$. It can be shown that $\zeta(s)$ extends to $\mathbb{C}$ as a meromorphic function with only a simple pole at $s=1$, with residue 11, and satisfies the functional equation $\xi(s)=\xi(1-s)$, where $\xi(z)=\frac{1}{2}z(z-1)\pi^{-z/2}\Gamma(\frac{1}{2}z)\zeta(z)$ and $\Gamma(w)=\int_{0}^{\infty}e^{-x}x^{w-1}\mathrm{d}x$. Let $\chi$ be a primitive Dirichlet character and $L(s,\chi)$ be the corresponding L-function. Most of the analytic properties for $\zeta$ are also known to hold for $L(s,\chi)$ [5, Chapter 10]. Let $\Theta_{\chi}$ be the supremum of the real parts of the zeros of $L(s,\chi)$. The Generalised Riemann Hypothesis (GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$. For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta,\sigma=\Re(s)$, $\gamma=\Im(\rho),\Lambda$ be the von-Mangoldt function and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
Lemma 1. (Theorem 4.11 of [7]). We have $\zeta(s)=\sum_{n\leq x}n^{-s}-\frac{x^{1-s}}{1-s}+O(x^{-\sigma})$ uniformly for $\sigma\geq\sigma_{0}>0$ and $|\Im(s)|<2\pi x/C$, where $C$ is a given constant >1>1.
@@
Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
@@
uniformly for $1\leq\sigma\leq 2$. By the residue theorem, we have
@@
where $0\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>0$ is a constant, $\chi(s)=\frac{\zeta(s)}{\zeta(1-s)}$ and $|\chi(\sigma-it)|=|\chi(\sigma+it)|=\Big|\Big(\frac{t}{2\pi}\Big)^{\frac{1}{2}-\sigma-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big|\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78]. Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1+it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
@@
for $1\leq\sigma\leq 2.$ Combining (1) and (6) gives $\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{8}T.$ By the log-derivative of the functional equation for $\zeta$, we have
@@
We now treat the ranges $m+1\leq 2n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq 2n\leq m-1$, where $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. For $m+1\leq 2n\leq m(1+\varepsilon)$, put $2n=m+r$ where $1\leq r\leq m\varepsilon-1$ thus $\log(2n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
@@
$=\frac{\Lambda(m)}{em^{iT}}\sum_{0<r\leq\varepsilon m-1}\Big(\frac{(m+r)^{iT}}{r}+O\Big(\frac{\varepsilon}{r\log m}\Big)\Big)$ (16)
@@
We now estimate
$J(T):=\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right).$
Let $\phi(r)=\frac{T}{2\pi}\log(1+\frac{r}{m})$ thus $\phi^{\prime}(r)=\frac{T}{2\pi(r+m)}.$ Let $\eta\in(0,1/2)$ be a constant, $\beta=\phi^{\prime}(T^{1/4})$ and $\alpha=\phi^{\prime}(m\varepsilon)$. By Lemma 4.10 of Titchmarsh [7, p.76], we have
$J(T)=\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi i(\phi(x)-vx)}\mathrm{d}x+O(T^{-1/4}\log T).$ (18)
Since $\beta-\alpha\ll\varepsilon\log^{2}T=o(1)$, note that the range $\alpha-\eta,\beta+\eta]$ contains at most one integer $v$. Hence (18) implies that
$J(T)\ll\Bigg|\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi i(\phi(x)-vx)}\mathrm{d}x\Bigg|+T^{-1/4}\log T.$ (19)
Define
$J_{0}(T)=\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi i(\phi(x)-vx)}\mathrm{d}x=\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi iT\Big[\log(1+\frac{x}{m})-\frac{vx}{T}\Big]}\mathrm{d}x.\\$ (20)
Let $G(x)=\frac{1}{x},F(x)=\log(1+\frac{x}{m})-\frac{vx}{T},\lambda=2\pi T$ and $F^{\prime}(x_{0})=0$ so that $\frac{T}{m+x_{0}}=v$ thus $F^{\prime\prime}(x_{0})=-\frac{v^{2}}{T}$ and $|T-mv|\sim\frac{Tx_{0}}{m}$ where $x_{0}\in[T^{1/4},m\varepsilon]$. Then $J_{0}(T)$ is an oscillatory integral of the form $\int_{a}^{b}G(x)\exp(i\lambda F(x))\mathrm{d}x$. Thus by the stationary phase approximation 222see e.g. https://grokipedia.com/page/Stationary phase approximation (the one-dimensional section)., we have
$J_{0}(T)\sim G(x_{0})e^{i\lambda F(x_{0})-i\pi/4}\sqrt{\frac{2\pi}{\lambda|F^{\prime\prime}(x_{0})|}}\ll G(x_{0})\sqrt{\frac{2\pi}{\lambda|F^{\prime\prime}(x_{0})|}}\ll\frac{1}{|T-mv|}\ll T^{-1/4}$ (21)
uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ as $\lambda=2\pi T\rightarrow\infty$. Combing (19), (20) and (21) yields
$\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right)\ll T^{-1/4}\log T$ (22)
uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. Applying (22) into (17) gives
$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log^{2}T).$ (23)
Putting $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O({Tr^{2}/m^{2}})}=1+O({Tr^{2}/m^{2}})$ into the right-hand side of (23) gives
$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{2}T).$ (24)
Summing (24) over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\ \leq T(1+\varepsilon)$ yields
@@
$+O(T^{3/4}\log^{2}T).$ (26)
For $\frac{m}{1+\varepsilon}\leq 2n\leq m-1,$ put 2n=mr2n=m-r where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ so that $\log(2n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$. Thus by arguing as above, one arrives at
@@
$=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log^{2}T).$ (30)
Let $\{{y\}}$ denote the fractional part of $y$ and $\mid\mid y\mid\mid$ denote the distance to the integer nearest to $y$. Note that if $u>0$ is not an integer, we have
[…]
as $X\rightarrow\infty$. 333see e.g. https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
[…]
Define
$U(T)=\sum_{V\leq m\leq T(1+\varepsilon)}\frac{1}{1+T^{1/4}\,\Bigl\|\frac{T}{2\pi m}\Bigr\|},\qquad V:=\frac{T}{(1+\varepsilon)\log^{2}T}.$
Set $Z=T^{1/4}$ and $x_{m}:=\frac{T}{2\pi m}.$ We split the sum $U(T)$ according to the size of $\|x_{m}\|$. For $\|x_{m}\|\leq Z^{-1}$, we have $\frac{1}{1+Z\|x_{m}\|}\asymp 1,$ and the corresponding contribution is bounded by
$\{m\in[V,T(1+\varepsilon)]:\|x_{m}\|\leq Z^{-1}\}.$
Fix $k\in\mathbb{N}$. The condition $\Bigl|\frac{T}{2\pi m}-k\Bigr|\leq Z^{-1}$ is equivalent to $\frac{T}{2\pi(k+Z^{-1})}\leq m\leq\frac{T}{2\pi(k-Z^{-1})}.$ The length of this interval is $\frac{T}{2\pi}\left(\frac{1}{k-Z^{-1}}-\frac{1}{k+Z^{-1}}\right)=\frac{T}{2\pi}\cdot\frac{2Z^{-1}}{k^{2}-Z^{-2}}\ll\frac{T}{Zk^{2}}$. Hence, for fixed $k$, the number of such integers $m$ is $\ll\frac{T}{Zk^{2}}+1.$ Since $x_{m}=\frac{T}{2\pi m}\in\Bigl[\frac{1}{2\pi(1+\varepsilon)},\frac{(1+\varepsilon)\log^{2}T}{2\pi}\Bigr],$ we have $1\leq k\ll(\log T)^{2}$. Summing over $k$ gives
$\{m:\|x_{m}\|\leq Z^{-1}\}\ll\sum_{k\leq(\log T)^{2}}\left(\frac{T}{Zk^{2}}+1\right)\ll\frac{T}{Z}.$
Thus the contribution from the $x_{m}$ for which $\|x_{m}\|\leq Z^{-1}$, is
$\sum_{\|x_{m}\|\leq Z^{-1}}\frac{1}{1+Z\|x_{m}\|}\ll\frac{T}{Z}=T^{3/4}.$ (33)
For $\|x_{m}\|>Z^{-1}$ we use $\frac{1}{1+Z\|x_{m}\|}\leq\frac{1}{Z\|x_{m}\|}.$ Partition $(0,1/2]$ dyadically: $2^{-j-1}<\|x_{m}\|\leq 2^{-j},0\leq j\leq\log_{2}Z.$ For such $m$, $\frac{1}{1+Z\|x_{m}\|}\ll\frac{2^{j}}{Z}$. Since $x_{m}$ is monotone in $m$, the number of $m$ in each dyadic interval is $\ll T\cdot 2^{-j}.$ Hence the contribution of each $j$ is $\ll T\cdot 2^{-j}\cdot\frac{2^{j}}{Z}=\frac{T}{Z}.$ Summing over $0\leq j\leq\log_{2}Z$ yields
$\sum_{\|x_{m}\|>Z^{-1}}\frac{1}{1+Z\|x_{m}\|}\ll\frac{T}{Z}\log Z\ll T^{3/4}\log T.$ (34)
Combining both ranges gives
$U(T)=\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{1}{1+T^{1/4}\Bigl\|\tfrac{T}{2\pi m}\Bigr\|}\ll T^{3/4}\log T.$ (35)
Inserting (32) into (30) and applying (35) gives
@@
$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big(\frac{1}{2}-\{{T/2\pi m\}}\Big)+O(T^{3/4}\log^{2}T).$ (37)
@@
since $\sum_{m\leq x}\Lambda(m)\sim x$ and $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.187]. Plugging (42) into (37) gives
@@
Remark 2. Our argument crucially relies on: (i) the functional equation for $\zeta$, and (ii) theorem 4.11 of [7]. Both (i) and (ii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers. Hence our approach wouldn’t work for such, whose analogues of the RH are known to be true. Also, (ii) does not hold for the Weil zeta functions, which are known to satisfy the RH.
@@
Remark 4. The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need our Lemma 1 to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log t)^{-1}$. However, this is not possible (as can be seen e.g by estimating $\zeta(a+it)=\sum_{n\leq\frac{T}{2}}n^{-a-it}+O(T^{-a})=\sum_{n\leq T/2}n^{(-a+\sigma_{t})-\sigma_{t}-it}+O(T^{-a})$ for any fixed $a>0$ via Abel summation, under the assumption that $\sum_{n\leq T/2}n^{-\sigma_{t}-it}=\zeta(\sigma_{t}+it)+O(1/\log t)$). It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}=o(1)$. Indeed, he begins his argument by quoting the Euler-McLaurin summation formula for $\zeta(\sigma+it)$ (equation 3.5.3). This is initially defined for $\sigma>1$, and then is extended by holomorphic continuation to $\sigma>0$. Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh assumes that $\sigma$ is bounded below by some constant $\sigma_{0}>0$. For the same reason, one cannot apply any of the best known zero-free regions of $\zeta$ (e.g. Vinogradov-Koborov) to bound $\Re(\rho_{T})$ in (58).